OpenStudy (anonymous):

How can I do this problem im not sure how to factor this Don't give me the answer just teach me please (4x - 5)(3x + 2) =0

OpenStudy (lalaly):

factor? its factored... or u want to find the values of x?

OpenStudy (anonymous):

i say quadtatic equation but thats just me

OpenStudy (anonymous):

Yeah the x value

OpenStudy (anonymous):

quadratic then

OpenStudy (anonymous):

Solve: (4x - 5)(3x + 2) =0

OpenStudy (lalaly):

u let 4x-5=0 and 3x+2=0 solve for x and u will have two values of x

OpenStudy (anonymous):

My teacher told me a different method that takes longer but is simpler?

OpenStudy (anonymous):

What is the official way

OpenStudy (anonymous):

O wrong question sorry